(x)^(2)+8x+(4y)^(2)-8y=80
Write the equation in standard from and then identify the graph of this equation and important characteristics of the graph, such as center, orientation, a,b,c.
*I am just so lost....*
Answer by stanbon(75887) 
 (x)^(2)+8x+(4y)^(2)-8y=80
Write the equation in standard from and then identify the graph of this equation and important characteristics of the graph, such as center, orientation, a,b,c.
==========================
Complete the square:
x^2+8x+? + 4(y^2-2y+#) = 80+? + 4#
------------------------------
x^2 + 8x + 16 + 4(y^2 - 2y + 1) = 80 + 16 + 4
(x+4)^2 + 4(y-1)^2 = 100
--------
(x+4)^2/100 + (y-1)^2/25 = 1
----
center: (-4,1)
orientation : wider than high
a = 10 ; b = 5 ; c = sqrt[100-25] = 5*sqrt(3)
================================================
Cheers,
Stan H.
=========================
| |
|
| |